Status Report

NASA Space Station On-Orbit Status 20 April 2005

By SpaceRef Editor
April 20, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 20 April 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  >>>Day 4 of joint Exp.10/Exp.11 operations. Day 189 in space for Exp. 10 (187 aboard ISS), with 5 days to go.

After wake-up at the regular time (2:00am EDT), both crews went to work on a busy schedule of ISS-10-to-ISS-11 handovers.

General handover activities continue to go well.  [As of yesterday evening, both expedition crews have completed several hours of dedicated handover activities and functional handover (joint walk-throughs on specific tasks), with Krikalev racking up 5h 50m functional, 2 h dedicated, and Phillips 4h 50m functional, 3h 40m dedicated.]

Before his breakfast, FE-10 Sharipov started on his first day of the Russian MBI-2 Diurez (“Diuresis”) experiment, for which he is the subject today and tomorrow.  [MBI-2 studies water-salt metabolism and volemic hormonal regulation during long duration spaceflight. Today, Salizhan was required to record toilet use time in the morning, perform urine sample collections throughout the day, and log meal and fluid intake on a record card. Additional work, to be completed tomorrow, consists of collecting his venous blood samples, followed by centrifugation to separate the blood into cells and plasma, and finally closeout and stowage.]

Afterwards, Sharipov underwent the Day 1 protocol of the Russian MedOps tests MBI-11 Gematologia (Hematology) which investigate the decrease in red blood cell mass during long duration exposure to micro-G, using the Reflotron-4 analyzer.  [CDR-11 Krikalev assisted in the test, drawing the blood samples for analysis with the special Erythrocyte kit after he and Salizhan had donned protective goggles.]


Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

After Sharipov had reconfigured the EGE-2 laptop and tagged up with ground specialists, he and Krikalev transferred & readied the new set of NEUROCOG experiment equipment, including the KOGNI kit and Halley instrument (new parts from Soyuz-216/10S delivery). Salizhan then had another extensive session with the experiment, focusing on virtual rotation in free floating and fixed position passages while recording EEG (electroencephalogram). [Assisted by CDR-11 Krikalev, the FE first activated the EGE-2 computer, then equipped himself with the Halley head electrodes. After doing the virtual turns/corridor episodes in fixed state (subject strapped down) and free-floating in zero-G called for by the Neurocog protocol, Sergei downloaded the EEG data to a PCMCIA memory card for subsequent return to Earth, dismantled the equipment and stowed it for future use.]

Sharipov also performed his third training session in the “Chibis” ODNT suit as part of his preparations for returning into gravity. Since there was no RGS/telemetry downlink, his vital body readings were obtained with the Tensoplus sphygmomanometer. A tagup with ground specialists via S-band supported the run, and Leroy Chiao assisted as required.  [The below-the-waist reduced-pressure device ODNT (US: LBNP) in the “Chibis” garment provides gravity-simulating stress to the body s cardiovascular/circulatory system for reestablishing the body s orthostatic tolerance (e.g., the Gauer-Henry reflex) after the six-month stay in zero-G. Gennady s and Mike s ODNT protocol today consisted of first imbibing 150-200 milliliters of water or juice, followed by a sequence of progressive regimes of reduced (negative) pressure, set at -25, -35, -40 and -45 mmHg, 5 minutes in each mode, while shifting from foot to foot at 10-12 steps per minute. The body s circulatory system interprets the pressure differential between upper and lower body as a gravity-like force pulling the blood (and other liquids) down.]

FE/SO-11 John Phillips supported the Huntsville POIC (Payloads Operations & Integration Center) in the checkout & annual re-certification of the European MSG (Microgravity Science Glovebox) by testing temperature and pressure sensors and working the MLC (MSG Laptop Computer) control panel for collecting data from the rack for comparison and analysis vis-à-vis previous test data.  [John received kudos from the MSG team for yesterday s work on the Glovebox. ( Everything powered up fine, operated well, and the software was successfully updated! )]

Continuing EVA-related activities in the US Airlock (A/L) started yesterday, Chiao and Phillips completed four hours of activities (broken up in smaller sequenced tasks) of initiating, then reconfiguring circulation/filtration of both EMU/spacesuit cooling loops, flushing with iodinated water, followed by swapping SCU-2 (Service & Cooling Umbilical 2) with an on-orbit spare, then checking and finally leak testing the new installation, along with the oxygen & water lines, communications, and power.  [The procedure utilized an EMU to circulate, filter, and flush out both EV-1 & EV-2 A/L cooling lines. The A/L EMU activities are serving the twofold purpose of being an important step on the Road to LF-1 and an excellent handover exercise. Following this activity, the A/L will again be available for EVA use.]

Using new 10S-delivered spares, John installed two LHAs (Lamp Housing Assemblies) and two BBAs (Baseplate Ballast Assembly) in the Node, one on port, the other on starboard (OP2-1,OS2-1). Two additional LHAs may have been installed in the Node later today if time permitted.  [Currently functioning lights: nine (of 12) in the Lab, four (of eight) in the Node, and three (of four) in the A/L.]

Phillips also conducted load troubleshooting on UOP-4 (Utility Outlet Panel #4) in the Lab, which had tripped RPC-4 (Remote Power Controller #4) last November, using the Aeolus volt/ohm scopemeter (for which a battery was recharged yesterday in the Pilobolus ).  [Recalling: Past work by Leroy Chiao on UOP4-associated hardware has exonerated a number of these loads as causes of the RPC trip.]

Both Russian air conditioners (SKV-1, SKV-2) are currently under investigation for decreased condensate production. The CCAA (Common Cabin Air Assembly) air conditioner in the US segment has picked up the load, filling up the Lab condensate collection tank more rapidly than expected. Leroy today hooked up a CWC (contingency water container) to the tank rack and transferred part of its contents.  [It is anticipated that another CWC will be needed for offloading the tank on/or about 4/23. US and RS experts are discussing water processing in the Russian segment (RS).]

Chiao conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various Russian segment (RS) hatchways, including the SM-to-Soyuz tunnel, and the FGB-to-Node passageway.  [This is especially important when the ventilation/circulation system has to cope with a large crew on board.]

FE-10 Sharipov completed the regular periodic download of data & imagery collected of the BIO-5 Rasteniya-2 (“Plants-2”) experiment to the computer for subsequent downlink to the ground. Today s activity also included recording plant harvesting with the DSR PD-150P video camera. The dry bean plants were bagged in silica gel and stowed for return to Earth.

VC8 guest cosmonaut Roberto Vittori continued busily performing experiments in his very crammed Eneide science program.  [Today, Vittori worked on NGF/Nerve Growth Factor (saliva collections), EST/Electronic Space Test (experiment activation, start time registration), CRISP-2/Crickets in Space (opening the KUBIK-AMBER lid for air exchange), HPA (Hand Posture Analysis (equipment setup in Lab by Sergei, ops, file transfer), LAZIO (hardware troubleshooting, experiment ops with mask, photo imagery), AGROSPACE/Beans-Seedlings (data checking & logging), ENE/Eneide (hardware deactivation, health check, setup and activation with an initial position/velocity vector for Session-1), MOP/Motion Perception (filling out questionnaire), and E-NOSE/Electronic Nose (hardware setup & activation for Session-1).]

The VC8 visiting crew also took extensive photo/video imagery of a number of Eneide experiments, in particular ASIA (Algorithm Study, Investigation & Analysis of electronic card sensitivity), E-NOSE, ETD (Eye Tracking Device), and FTS (Italian Food Tasting in Space).

Krikalev and Sharipov, later joined by Vittori, had time reserved for scheduled commemorative (Russian: symbolic ) activity, as is standard tradition for visiting guest crewmembers. It usually consists of signing and stamping envelopes and imaging other memorabilia.   [VC8 items to be returned by the Expedition 10 crew are 60 envelopes with the VC8 mission logo, a Microcosmos kit with a clear crystal with an embedded nanotechnology silicon chip bearing a picture of Mona Lisa by Leonardo da Vinci in Italy (500 years anniversary in 2005) painted by Russian artist Georgy Pusenkov with pixels, reduced by nanotechnology (courtesy a Zelenograd company) to individual atoms, and the Dove of Peace kit for photography, to be demonstrated during a future downlink conference.]

The E10 crew conducted a slightly abbreviated physical exercise program on TVIS treadmill, RED exerciser, and VELO cycle with bungee cord load trainer. Leroy then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

Salizhan did the daily routine maintenance of the SM’s SOZh environment control & life support system, including its ASU toilet system and today also the weekly inspection of the BRPK air/liquid condensate separator apparatus. In addition, he prepared the regular IMS (inventory management system) delta file for export/import to the IMS databases.

Using the automatic temperature recorder (ART), Sharipov performed the regular temperature check on the BIO-11 Statokonia payload with the ULITKA ( snail ) incubator, set up in the SM since 3/3.  [BIO-11 studies the composition of statoconia, i.e., the organ of equilibrium in snails, and other phenomena exhibited by ulitka in zero-G and post-flight.]

At 8:30am EDT, Roberto Vittori set up the SM’s amateur radio equipment (Ericsson VHF transceiver, headset, power supply) for another ARISS session, and at 8:35am he engaged in a 10-min. ham radio exchange with students at ESA s ESRIN Center at Frascatti, Italy, who represented the winning classes of a European competition for primary schools in Italy s Lazio region. As student mementos, Sharipov took photographs of the activity with the Nikon D1X before and during the ham radio pass.

After rave reviews for yesterday s TV interviews by ABC & CNN for the five-member crew, at 1:20pm today they conducted another live TV PAO exchange with US media (Fox News/Linda Vester, CBS News/Bill Harwood, Peter King).

At 1:30pm, FE-VC8 Vittori conducted his daily 10-min tagup with ESA Ops personnel via audio/S-band.

Later, at 2:15pm, Roberto supported another live TV downlink with the Italian TV channel TG3 Ragazzi, RAI (Paola Sensini) in Rome, with Sergei Krikalev acting as camera operator.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links


Working off his discretionary job jar task list, Roberto Vittori again used his Nikon D1X digital camera for shooting pictures of his home country, for which two overflight opportunities were uplinked.

The Elektron was shut down yesterday evening (3:38pm EDT) due to increasing voltage in the BZh electrolyzer. TsUP/Moscow is developing a new plan. As of last night, oxygen partial pressure (ppO2) was good at 165.3 mmHg/Torr (lower FR limit is 152 mmHg, incl. error band).

Researchers on the ground again conducted the European/Russian laser-beam experiment SPQR (Specular Point-like Quick Reference), today during a 5-min window at 8:42am.  [SPQR, installed at SM window #3 along with its Nikon D1X camera, tests a ground-based imaging system, using special optics and image processing, to determine the feasibility of detecting external damage to a spacecraft in orbit from the ground. It uses a pyramidal corner reflector (CCR, Cube Corner Reflector) at the SM window, to reflect a laser beam emitted by a ground station back to the ground. The crew was advised not to look out the portholes during the sessions, the times of which were uplinked, and there are no CEO targets scheduled during the brief sessions. The SPQR Hazard Report indicates that the laser power at the ISS remains well below the threshold for injury.]

ISS Orbit  (as of this morning, 8:32am EDT [= epoch]):

  • Mean altitude — 355.0 km
  • Apogee height — 359.9 km
  • Perigee height — 350.1 km
  • Period — 91.64 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0007248
  • Solar Beta Angle — 15.1 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.71
  • Mean altitude loss in last 24 hours — 95 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 36661

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

Expedition 10 Flight Timelines:

Soyuz 9S (Expedition 10+1; Leroy Chiao, Salizhan Sharipov, Roberto Vittori):

  • Undocking from FGB — 4/24 (Sun.), 2:44pm EDT;
  • Sep Burn #1 (manual) — 2:50pm;
  • Deorbiting Burn — 5:24pm (4 min 23 sec, delta-V 115.2 m/s);
  • Module Sep —  5:49pm;
  • Atmospheric Entry — 5:50pm;
  • Landing in darkness — 4/24 (Sun.) 6:09pm EDT; 3:09am (4/25) local Kustanai/Kazakhstan;
  • Sunrise at Kustanai landing site — 5:16am. [Note: Kazakhstan remains on Standard Time; thus: local time = GMT+5].

Return to Flight:

  • LF-1 (STS-114)/Increment 11 SORR (Stage Operations Readiness Review) — 4/22 (Fri) at JSC;
  • LF1 (STS-114)/Discovery launch windows (all times EDT), for FD3 docking:
  • 5/15: 3:45 – 3:55pm;
  • 5/16: 2:22 – 2:32pm;
  • 5/17: 1:59 – 2:07pm;
  • 5/18: 1:34 – 1:44pm;
  • 5/19: 1:12 – 1:22pm;
  • 5/20: 12:47 – 12:57pm;
  • 5/21: 12:24 – 12:34pm;
  • 5/22: 11:58 – 12:08pm;
  • 5/23: 11:36 – 11:46pm;
  • _ etc.

Note: For the May/June launch period, the daily 10-minute planar launch window (i.e., in ISS orbit plane) starts an average 23 minutes earlier each day, extends into early June and closes due to current constraints of Daylight Launch (6/7) or ET umbilical photo opportunity (6/3). Figures are approximate. There are additional opportunities for docking on FD4 (Flight Day 4), not planned. If STS-114 launches on 5/15, docking will be on 5/17 and undocking on 5/25.

Other Upcoming Main Events:

  • Progress M-53 (18P) launch 6/10;
  • ULF1.1 (STS-121) launch — NET 7/12;
  • Progress M-54 (19P) launch — 8/24;
  • Soyuz TMA-7 (11S) launch — 9/27.

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.